# VQ-Style: Disentangling Style and Content in Motion with Residual Quantized Representations

### v1

@@ Abstract
Human motion data is inherently rich and complex, containing both semantic content and subtle stylistic features that are challenging to model. We propose a novel method for effective disentanglement of the style and content in human motion data to facilitate style transfer. Our approach is guided by the insight that content corresponds to coarse motion attributes while style captures the finer, expressive details. To model this hierarchy, we employ Residual Vector Quantized Variational Autoencoders (RVQ-VAEs) to learn a coarse-to-fine representation of motion. We further enhance the disentanglement by integrating contrastive learning and a novel information leakage loss with codebook learning to organize the content and the style across different codebooks. We harness this disentangled representation using our simple and effective inference-time technique Quantized Code Swapping, which enables motion style transfer without requiring any fine-tuning for unseen styles. Our framework demonstrates strong versatility across multiple inference applications, including style transfer, style removal, and motion blending.
@@ 1 Introduction
Virtual characters play a central role in media and entertainment, from animation and video games to immersive experiences. However, creating realistic and expressive character animation remains a labor-intensive process that often requires extensive manual work from artists, making it both time-consuming and costly. Recent research in character animation has explored how modern machine learning techniques can reduce repetitive work and automate aspects of the animation pipeline. One notable problem is style transfer, where the style of a motion—such as walking happily or angrily—is transferred from one motion clip to another, while preserving the content, or semantic meaning, of the latter motion; usually referred to as the content clip.
Achieving compelling motion style transfer requires disentangling style from content—a fundamental and challenging problem in motion representation. While humans naturally exhibit a wide range of walking styles, clearly defining and separating style and content in motion data is non-trivial. Our goal is to learn a representation that effectively disentangles the content and style from motion clips and allows for transferring styles realistically. We approach this problem by interpreting content as the coarse and structural attributes of motion, and style as the finer details that introduce expressive nuances to the movement. To achieve this separation, we employ Residual Vector Quantized Variational Autoencoders (RVQ-VAEs) [LKK22] to learn a coarse-to-fine representation. This is complemented by novel contrastive and mutual information losses that prevent style leakage into content. We demonstrate that the resulting representation effectively separates style from content and enables style transfer through Quantized Code Swapping. Our approach performs style transfer as an inference-only task, requiring no additional fine-tuning for unseen styles. Furthermore, unlike prior works, our model is trained without adversarial or cyclic training, enabling a more stable convergence.
We evaluate our method on multiple motion capture datasets [MSK22, AWL20, XWCH15] and show that our representation embeds the content in the initial codebooks and the style in latter ones. Our experiments show effective style transfer across both seen and unseen stylistic motions, measured via style classifier accuracy, while maintaining the original content, measured via trajectory deviation. In addition, it enables further capabilities including smooth transitions between styles, content extraction, motion interpolation, and content-style mixing for data augmentation.
@@
  • Our framework supports several applications at inference—including style transfer and style transitions over arbitrary-length motions—via swapping and blending operations on the disentangled quantized codebooks.
@@ 2 Related Work
[XWCH15] were one of the first to apply style transfer to motion data using a motion database and a KNN search. They first pre-process the database to track the closest neighboring pose for each pose in the dataset. Afterwards, they can perform real-time style transfer relying on a KNN search and simple linear transformation. In the context of deep learning, [AWL20] learn separate style and temporal content codes. They use Instance Normalization (IN) and Adaptive Instance Normalization (AdaIN) to remove the original style from the content motion and reintroduce the desired style codes, respectively. Since the model relied heavily on style labels and a discriminator, it struggles to generalize to unseen styles. Both these works also released their annotated styled motion data that have been used by many works since.
@@
[PJL21] use a graph convolutional network (GCN) where the pooling and unpooling operations are performed according to the human skeleton. In addition, they learn a mapping network from random noise to a style code to perform style transfer from unseen styles at inference time. Motion Puzzle [JPL22] also uses a similar GCN, but trains from one source motion to multiple target motions each to stylize one or more body parts. This lets them combine different styles across body parts during inference. More recently, [TWW24] used contact information for style transfer via learning separate GCNs to extract contact, trajectory and style information from the source motion. They can then perform style transfer at inference by swapping the style codes passed to the generator transformer. Similarly, the contact timing can also be swapped between the style and content motion.
@@
100STYLE dataset[MSK22] is a locomotion dataset with a combination of 100 different styles and different gaits (walking, running, & strafing). In addition to introducing the dataset, [MSK22] train a mixture of expert (MoE) model with a FiLM [DPS18, PSDV18] generator and a pre-trained periodic autoencoder [SMK22] providing learned phase labels. Although their approach works well for seen styles, they require fine-tuning to adapt to unseen styles. [TWW23] also use an MoE model along with the periodic autoencoder to learn a motion VAE to reconstruct the next frame conditioned on the current frame. The decoder can be conditioned on a style token to perform style transfer. [DWF25] further extend the periodic autoencoder to learn body part-wise phase labels. They can then encode these part-wise phase labels into style labels, allowing mixing styles across body parts similar to Motion Puzzle. These methods rely on a pretrained phase manifold for motion, a MoE and a motion sampler, constructing a complex training pipeline.
@@
Diffusion models have also been employed for motion style transfer. [RGS24] show that for a pre-trained motion diffusion model [TRG23], the queries and keys in the self-attention blocks correspond to motion outline and motifs respectively. Using noise inversion and swapping the queries and keys between a content and a style motion clips, they can perform zero-shot style transfer. In contrast, [ZXJ24] learn a ControlNet [ZRA23] based style adapter and combine classifier-based and classifier-free [HS22] guidance to adapt a pre-trained text-to-motion model for stylization. [SJL24] specifically train a diffusion model for style transfer and first learn a multi-condition extractor that can extract trajectory and remove style from the content motion, and extract the style from the style motion. A latent diffusion model is then trained on these features to stylize the content motion.
@@
While the above-mentioned diffusion models offer promising results, their iterative generative process makes them too slow for real-time applications and for stylizing motion of arbitrary lengths. Hence, [GMZ24] first learn a motion latent representation using a VAE [KW19]. This representation is then disentangled using style and content encoders and a generator. Reconstruction, homo-style alignment, and cycle consistency losses are used to ensure that the new latent space is sufficiently disentangled. Although they can transfer unseen styles, their method requires training multiple models, and losses such as homo-style alignment and cycle consistency can make the training more unstable.
@@
To our knowledge, we are the first work to explore VQ-VAEs [VDOV17] in the context of motion stylization. We show that using a Residual VQ-VAE (RVQ-VAE) [LKK22], we can learn a disentangled latent space for motion style transfer without specialized style and content encoders and perform style transfer for arbitrary long motion sequences in real time.
@@ 3 Method
Figure 1: Our approach encodes motion into a number of codebooks stacked in a residual manner. Our training strategy enables the content to be represented by the first (blue) codebook, while style is represented by the downstream codebooks in red.
@@
Our goal is to learn a motion representation that effectively disentangles the stylistic characteristics of human motion, from its semantic content. We utilize an RVQ-VAE [YSZ24, LKK22] framework that consists of an encoder $\mathcal{E}$, a decoder $\mathcal{D}$ and the latent space is modeled by multiple categorical distributions or codebooks $\mathcal{B}$. Overview of our framework is demonstrated in Fig. 1. We propose a training strategy in which the initial codebooks capture the content of the motion, while the subsequent codebooks encode the stylistic nuances. Key components of our training strategy include contrastive learning (Sec. 3.3) and mutual information loss (Sec. 3.4). This disentangled representation can be utilized at inference time to perform style transfer via Quantized Code Swapping (Sec. 3.5).
@@ 3.1 Quantized Motion Embedding
RVQ-VAEs can be trained by sampling a random number of residual codebooks in use, e.g. for each training sample, only the first $n$ number of codebooks $\mathcal{B}_{0},\mathcal{B}_{1},...,\mathcal{B}_{n-1}]$ are employed to reconstruct the data. This training strategy effectively enables learning a coarse-to-fine representation [YSZ24, LKK22], where the initial codebooks encode coarse information and the subsequent codebooks capture finer details. Intuitively, the content of a motion clip corresponds to these coarse components of the data, while stylistic variations are expressed through finer nuances. Under this perspective, RVQ-VAEs provide a natural and effective framework for disentangling style from content in motion.
@@ 3.2 Training
To effectively train the codebooks $\mathcal{B}_{i}$, we use commitment loss and Exponential Moving Average (EMA) update as detailed in Appendix A. Code reset is also done similar to [ŁCS20] for codes that are unused during training to encourage more uniform usage of all codes in the codebook.
@@ 3.3 Contrastive Learning
For this purpose, we adopt the Multi-Pos contrastive loss introduced in [TFI23] to contrast all pairs of positive and negative samples within a batch. Let $\mathbf{a}$ denote an anchor embedding sampled from the batch, and $\{\mathbf{b}_{1},\mathbf{b}_{2},...,\mathbf{b}_{x}\}$ denote the remainder. The Multi-Pos contrastive loss is given by the cross entropy, $H$, between similarity-based distribution of how closely $\mathbf{a}$ matches each $\mathbf{b}$ and target distribution constructed from ground truth style labels.
@@
[HC20] introduced contrastive learning within a VQ-VAE framework, simply by applying the contrastive objective to the continuous embeddings prior to quantization. In contrast, our method applies contrastive loss directly to the residual embeddings, i.e. after quantization. We show that this choice enables backpropagation to update the codes in the latest codebook, without explicitly affecting the gradients of the earlier stages in the pipeline. This is particularly crucial for our goal, since the style-related contrasts should not affect the initial codebooks which are intended to encode content. To show this, note that gradients of the forward path are computed using straight-through method [HCAI23]:
@@
A second key distinction from [HC20] lies in how the codebooks are optimized. While their method uses the standard VQ objective, we adopt the more robust EMA update (see Appendix A). Because EMA updates are applied manually rather than through backpropagation, they break differentiability. Therefore, incorporating the contrastive loss into codebook learning requires careful design. Although the contrastive loss can update the codebook via backpropagation, EMA is typically performed in the forward step and prior to backpropagation, invalidating the gradients. To properly incorporate the contrastive updates, the EMA update must instead be applied after backpropagation, as a second update step.
@@ 3.4 Mutual Information Loss
The contrastive loss on style labels enables disentanglement of different styles in the residual codebooks. However, style information can still be stored in the content codebook. Incorporating a similar content contrastive loss for the content codebooks is infeasible as semantic rich content labels are not always available. Instead, we propose restricting the model from inferring style labels from the content codebooks. This can be achieved using a mutual information loss [FH61, NYC19]. Unlike prior work that typically maximize mutual information between given labels and embeddings, we minimize the mutual information between the selected content codes, $z$, and style labels, $l$. This ensures that no style can be inferred from the content embeddings. This is formulated as,
@@ 3.5 Quantized Code Swapping
Figure 2: Quantized Latent Code Swapping. To transfer style from one clip onto another, we encode both clips, and decode a combined embedding constructed by adding the content code (shown in blue) from the content clip and the style code (shown in red) from the second clip.
@@ 4.1 Style-Content Disentanglement
Low-dimensional projection. Fig. 3 presents a TSNE [VdMH08] visualization of the residual embeddings $\mathbf{r}_{i}$ for each layer, color-coded by the style labels. As shown in Fig. 3(a), even without contrastive learning, the RVQ-VAE architecture already clusters motions of same style close together in the second codebook. When the contrastive loss is introduced, the separation between styles becomes more distinct (Fig. 3(b)). Adding the mutual information loss further strengthens the disentanglement (Fig. 3(c)). Notably, the model also exhibits disentanglement for styles that were unseen during training (Fig. 3(d)).
@@ 4.6 Data Augmentation
As demonstrated in previous works [AKB24, MU22], data augmentation and bias reduction can significantly enhance motion generation with deep learning models. Our interpretable latent space also enables new data augmentation methods.
@@ 4.7 Quantitative analysis
Table 1: Baseline comparison on 100STYLE dataset with LPN-Style[MSK22]. Numbers in parenthesis show top-5 accuracy. LPN-Style denotes the experiment excluding 8 broken styles (those with accuracy lower than 10%).
@@
Table 2: Comparison to [GMZ24] on Aberman and Xia datasets. Numbers in parenthesis show top-3 accuracy.
@@
In this section, we quantitatively compare our method with the baselines [MSK22] and [GMZ24], in terms of style preservation accuracy and content trajectory deviation. We compare against the motion-based supervised setting for [GMZ24] as it is closest to our inference configuration. We define style preservation accuracy, $A_{S}$ as the classification accuracy of a pretrained style classifier over the generated motion. Content trajectory deviation, $D_{C}$ measures the deviation of the generated motion’s root trajectory from that of the content motion. For more details on the metric definition and dataset splits, please refer to Appendix C.
@@ 4.7.1 Style Accuracy
In Table 1, we report style accuracy, $A_{S}$, for the test and the unseen style subsets of the 100STYLE [MSK22] dataset. As a baseline, we use the locomotion controller LPN-Style provided by [MSK22] to generate stylized motion to measure their test $A_{S}$. In practice, we observed that LPN-Style produced severely broken or incorrectly stylized motion for certain styles, with accuracies falling below 10%. To ensure a fair evaluation, we report their performance both including and excluding these problematic styles in the test set. Our method scores top-1 $A_{S}$ of 83.20% on the test set, outperforming both versions of LPN-Style.
For unseen styles, our method achieves 68.95% top-1 and 98.83% top-5 $A_{S}$, whereas LPN-Style is incapable of zero-shot generalization. Since LPN-Style cannot generate motion for unseen styles, they must fine-tune their FiLM module [PSDV18] on the new styles. This fine-tuning improves their performance to 86.92% $A_{S}$, while our similarly fine-tuned model can achieve a score of 96.88%, significantly outperforming LPN-Style.
We evaluate our method on the Aberman [AWL20] and Xia [XWCH15] datasets which contain more generic stylized motions beyond locomotion. For training, we used the retargeted motions from Aberman provided by GenMoStyle[GMZ24]. Our model is trained solely on the Aberman training set, but we evaluate it both on the Aberman and Xia datasets. Since these datasets contain fewer styles (16 for Aberman and 8 for Xia), we report top-3 $A_{S}$ rather than top-5 $A_{S}$ used for 100STYLE.
As reported in Table 2, our method outperforms GenMoStyle for style accuracy across all subsets and both top-1 and top-3 $A_{S}$. Notably, while GenMoStyle do not use the Xia dataset to train their style and content encoders, their general motion autoencoder is pretrained on all CMU[CMU], Aberman and Xia datasets. In contrast, we use the Xia dataset exclusively for evaluation.

### v2

@@ Abstract
Human motion data is inherently rich and complex, containing both semantic content and subtle stylistic features that are challenging to model. We propose a novel method for effective disentanglement of the style and content in human motion data to facilitate style transfer. Our approach is guided by the insight that content corresponds to coarse motion attributes while style captures the finer, expressive details. To model this hierarchy, we employ Residual Vector Quantized Variational Autoencoders (RVQ-VAEs) to learn a coarse-to-fine representation of motion. We further enhance the disentanglement by integrating codebook learning with contrastive learning and a novel information leakage loss to organize the content and the style across different codebooks. We harness this disentangled representation using our simple and effective inference-time technique Quantized Code Swapping, which enables motion style transfer without requiring any fine-tuning for unseen styles. Our framework demonstrates strong versatility across multiple inference applications, including style transfer, style removal, and motion blending.
@@ 1 Introduction
Virtual characters play a central role in media and entertainment, from animation and video games to immersive experiences. However, creating realistic and expressive character animation remains a labor-intensive process that often requires extensive manual work from artists, making it both time-consuming and costly. Recent research in character animation has explored how modern machine learning techniques can reduce repetitive work and automate aspects of the animation pipeline. One notable problem is style transfer, where the style of a motion—such as walking happily or angrily—is transferred from one motion clip to another, while preserving the content, or semantic meaning, of the latter motion, which is usually referred to as the content clip.
Achieving compelling motion style transfer requires disentangling style from content—a fundamental and challenging problem in motion representation. While humans naturally exhibit a wide range of walking styles, clearly defining and separating style and content in motion data is non-trivial. Our goal is to learn a representation that effectively disentangles the content and style from motion clips and allows for transferring styles realistically. We approach this problem by interpreting content as the coarse and structural attributes of motion, and style as the finer details that introduce expressive nuances to the movement. To achieve this separation, we employ Residual Vector Quantized Variational Autoencoders (RVQ-VAEs) [16] to learn a coarse-to-fine representation. This is complemented by novel contrastive and mutual information losses that prevent style leakage into content. We demonstrate that the resulting representation effectively separates style from content and enables style transfer through Quantized Code Swapping. Our approach performs style transfer as an inference-only task, requiring no additional fine-tuning for unseen styles. Furthermore, unlike prior works, our model is trained without adversarial or cyclic training, enabling a more stable convergence.
We evaluate our method on multiple motion capture datasets [18, 1, 33] and show that our representation embeds the content in the initial codebooks and the style in latter ones. Our experiments show effective style transfer across both seen and unseen stylistic motions, measured via style classifier accuracy, while maintaining the original content, measured via trajectory deviation. In addition, it enables further capabilities including smooth transitions between styles, content extraction, motion interpolation, and content-style mixing for data augmentation.
@@
  • Our framework supports several applications at inference, including style transfer and style transitions over arbitrary-length motions, via swapping and blending operations on the disentangled quantized codebooks.
@@ 2 Related Work
[33] were one of the first to apply style transfer to motion data using a motion database and a KNN search. They first pre-process the database to track the closest neighboring pose for each pose in the dataset. Afterwards, they can perform real-time style transfer relying on a KNN search and simple linear transformation. In the context of deep learning, [1] learn separate style and temporal content codes. They use Instance Normalization (IN) and Adaptive Instance Normalization (AdaIN) to remove the original style from the content motion and reintroduce the desired style codes, respectively. Since the model relied heavily on style labels and a discriminator, it struggles to generalize to unseen styles. Both these works also released their annotated styled motion data that have been used by many works since.
@@
[20] use a graph convolutional network (GCN) where the pooling and unpooling operations are performed according to the human skeleton. In addition, they learn a mapping network from random noise to a style code to perform style transfer from unseen styles at inference time. Motion Puzzle [13] also uses a similar GCN, but trains from one source motion to multiple target motions each to stylize one or more body parts. This lets them combine different styles across body parts during inference. More recently, [27] used contact information for style transfer via learning separate GCNs to extract contact, trajectory and style information from the source motion. They can then perform style transfer at inference by swapping the style codes passed to the generator transformer. Similarly, the contact timing can also be swapped between the style and content motion.
@@
100STYLE dataset[18] is a locomotion dataset with a combination of 100 different styles and different gaits (walking, running, & strafing). In addition to introducing the dataset, [18] train a mixture of expert (MoE) model with a FiLM [5, 21] generator and a pre-trained periodic autoencoder [25] providing learned phase labels. Although their approach works well for seen styles, they require fine-tuning to adapt to unseen styles. [26] also use an MoE model along with the periodic autoencoder to learn a motion VAE to reconstruct the next frame conditioned on the current frame. The decoder can be conditioned on a style token to perform style transfer. [4] further extend the periodic autoencoder to learn body part-wise phase labels. They can then encode these part-wise phase labels into style labels, allowing mixing styles across body parts similar to Motion Puzzle. These methods rely on a pretrained phase manifold for motion, a MoE and a motion sampler, constructing a complex training pipeline.
@@
Diffusion models have also been employed for motion style transfer. [22] show that for a pre-trained motion diffusion model [29], the queries and keys in the self-attention blocks correspond to motion outline and motifs respectively. Using noise inversion and swapping the queries and keys between a content and a style motion clips, they can perform zero-shot style transfer. In contrast, [36] learn a ControlNet [35] based style adapter and combine classifier-based and classifier-free [10] guidance to adapt a pre-trained text-to-motion model for stylization. [24] specifically train a diffusion model for style transfer and first learn a multi-condition extractor that can extract trajectory and remove style from the content motion, and extract the style from the style motion. A latent diffusion model is then trained on these features to stylize the content motion.
@@
While the above-mentioned diffusion models offer promising results, their iterative generative process makes them too slow for real-time applications and for stylizing motion of arbitrary lengths. Hence, [7] first learn a motion latent representation using a VAE [14]. This representation is then disentangled using style and content encoders and a generator. Reconstruction, homo-style alignment, and cycle consistency losses are used to ensure that the new latent space is sufficiently disentangled. Although they can transfer unseen styles, their method requires training multiple models, and losses such as homo-style alignment and cycle consistency can make the training more unstable.
@@
To our knowledge, we are the first work to explore VQ-VAEs [31] in the context of motion stylization. We show that using a Residual VQ-VAE (RVQ-VAE) [16], we can learn a disentangled latent space for motion style transfer without specialized style and content encoders and perform style transfer for arbitrary long motion sequences in real time.
@@ 3 Method
Figure 1: Our approach encodes motion into a number of codebooks stacked in a residual manner. Our training strategy enables the content to be represented by the first (blue) codebook, while style is represented by the downstream codebooks (in red).
@@
Our goal is to learn a motion representation that effectively disentangles the stylistic characteristics of human motion, from its semantic content. We utilize an RVQ-VAE [34, 16] framework that consists of an encoder $\mathcal{E}$, a decoder $\mathcal{D}$ and the latent space is modeled by multiple categorical distributions or codebooks $\mathcal{B}$. Overview of our framework is demonstrated in Fig. 1. We propose a training strategy in which the initial codebooks capture the content of the motion, while the subsequent codebooks encode the stylistic nuances. Key components of our training strategy include contrastive learning (Sec. 3.3) and mutual information loss (Sec. 3.4). This disentangled representation can be utilized at inference time to perform style transfer via Quantized Code Swapping (Sec. 3.5).
@@ 3.1 Quantized Motion Embedding
RVQ-VAEs can be trained by sampling a random number of residual codebooks in use, e.g. for each training sample, only the first $n$ number of codebooks $\mathcal{B}_{0},\mathcal{B}_{1},...,\mathcal{B}_{n-1}]$ are employed to reconstruct the data. This training strategy effectively enables learning a coarse-to-fine representation [34, 16], where the initial codebooks encode coarse information and the subsequent codebooks capture finer details. Intuitively, the content of a motion clip corresponds to these coarse components of the data, while stylistic variations are expressed through finer nuances. Under this perspective, RVQ-VAEs provide a natural and effective framework for disentangling style from content in motion.
@@ 3.2 Training
To effectively train the codebooks $\mathcal{B}_{i}$, we use commitment loss and Exponential Moving Average (EMA) update as detailed in Appendix A. Code reset is also done similar to [15] for codes that are unused during training to encourage more uniform usage of all codes in the codebook.
@@ 3.3 Contrastive Learning
For this purpose, we adopt the Multi-Pos contrastive loss introduced in [30] to contrast all pairs of positive and negative samples within a batch. Let $\mathbf{a}$ denote an anchor embedding sampled from the batch, and $\{\mathbf{b}_{1},\mathbf{b}_{2},...,\mathbf{b}_{x}\}$ denote the remainder. The Multi-Pos contrastive loss is given by the cross entropy, $H$, between similarity-based distribution of how closely $\mathbf{a}$ matches each $\mathbf{b}$ and target distribution constructed from ground truth style labels.
@@
[9] introduced contrastive learning within a VQ-VAE framework, simply by applying the contrastive objective to the continuous embeddings prior to quantization. In contrast, our method applies contrastive loss directly to the residual embeddings, i.e. after quantization. We show that this choice enables backpropagation to update the codes in the latest codebook, without explicitly affecting the gradients of the earlier stages in the pipeline. This is particularly crucial for our goal, since the style-related contrasts should not affect the initial codebooks which are intended to encode content. To show this, note that gradients of the forward path are computed using straight-through method [12]:
@@
A second key distinction from [9] lies in how the codebooks are optimized. While their method uses the standard VQ objective, we adopt the more robust EMA update (see Appendix A). Because EMA updates are applied manually rather than through backpropagation, they break differentiability. Therefore, incorporating the contrastive loss into codebook learning requires careful design. Although the contrastive loss can update the codebook via backpropagation, EMA is typically performed in the forward step and prior to backpropagation, invalidating the gradients. To properly incorporate the contrastive updates, the EMA update must instead be applied after backpropagation, as a second update step.
@@ 3.4 Mutual Information Loss
The contrastive loss on style labels enables disentanglement of different styles in the residual codebooks. However, style information can still be stored in the content codebook. Incorporating a similar content contrastive loss for the content codebooks is infeasible as semantic rich content labels are not always available. Instead, we propose restricting the model from inferring style labels from the content codebooks. This can be achieved using a mutual information loss [6, 19]. Unlike prior work that typically maximize mutual information between given labels and embeddings, we minimize the mutual information between the selected content codes, $z$, and style labels, $l$. This ensures that no style can be inferred from the content embeddings. This is formulated as,
@@
Figure 2: Quantized Latent Code Swapping. To transfer style from one clip onto another, we encode both clips, and decode a combined embedding constructed by adding the content code (blue) from the content clip and the style code (red) from the second clip.
@@ 4.1 Style-Content Disentanglement
Low-dimensional projection. Fig. 3 presents a TSNE [32] visualization of the residual embeddings $\mathbf{r}_{i}$ for each layer, color-coded by the style labels. As shown in Fig. 3(a), even without contrastive learning, the RVQ-VAE architecture already clusters motions of same style close together in the second codebook. When the contrastive loss is introduced, the separation between styles becomes more distinct (Fig. 3(b)). Adding the mutual information loss further strengthens the disentanglement (Fig. 3(c)). Notably, the model also exhibits disentanglement for styles that were unseen during training (Fig. 3(d)).
@@ 4.6 Data Augmentation
As demonstrated in previous works [2, 17], data augmentation and bias reduction can significantly enhance motion generation with deep learning models. Our interpretable latent space also enables new data augmentation methods.
@@ 4.7 Quantitative analysis
Table 1: Baseline comparison on 100STYLE dataset with LPN-Style[18]. Numbers in parenthesis show top-5 accuracy. LPN-Style denotes the experiment excluding 8 broken styles (those with accuracy lower than 10%).
@@
Table 2: Comparison to [7] on Aberman and Xia datasets. Numbers in parenthesis show top-3 accuracy.
@@
In this section, we quantitatively compare our method with the baselines [18] and [7], in terms of style preservation accuracy and content trajectory deviation. We compare against the motion-based supervised setting for [7] as it is closest to our inference configuration. We define style preservation accuracy, $A_{S}$ as the classification accuracy of a pretrained style classifier over the generated motion. Content trajectory deviation, $D_{C}$ measures the deviation of the generated motion’s root trajectory from that of the content motion. For more details on the metric definition and dataset splits, please refer to Appendix C.
@@ 4.7.1 Style Accuracy
In Table 1, we report style accuracy, $A_{S}$, for the test and the unseen style subsets of the 100STYLE [18] dataset. As a baseline, we use the locomotion controller LPN-Style provided by [18] to generate stylized motion to measure their test $A_{S}$. In practice, we observed that LPN-Style produced severely broken or incorrectly stylized motion for certain styles, with accuracies falling below 10%. To ensure a fair evaluation, we report their performance both including and excluding these problematic styles in the test set. Our method scores top-1 $A_{S}$ of 83.20% on the test set, outperforming both versions of LPN-Style.
For unseen styles, our method achieves 68.95% top-1 and 98.83% top-5 $A_{S}$, whereas LPN-Style is incapable of zero-shot generalization. Since LPN-Style cannot generate motion for unseen styles, they must fine-tune their FiLM module [21] on the new styles. This fine-tuning improves their performance to 86.92% $A_{S}$, while our similarly fine-tuned model can achieve a score of 96.88%, significantly outperforming LPN-Style.
We evaluate our method on the Aberman [1] and Xia [33] datasets which contain more generic stylized motions beyond locomotion. For training, we used the retargeted motions from Aberman provided by GenMoStyle[7]. Our model is trained solely on the Aberman training set, but we evaluate it both on the Aberman and Xia datasets. Since these datasets contain fewer styles (16 for Aberman and 8 for Xia), we report top-3 $A_{S}$ rather than top-5 $A_{S}$ used for 100STYLE.
As reported in Table 2, our method outperforms GenMoStyle for style accuracy across all subsets and both top-1 and top-3 $A_{S}$. Notably, while GenMoStyle do not use the Xia dataset to train their style and content encoders, their general motion autoencoder is pretrained on all CMU[3], Aberman and Xia datasets. In contrast, we use the Xia dataset exclusively for evaluation.